OpenStudy (anonymous):

need help plz: simplify:8 times the square root of 63 x to the fifth power

OpenStudy (anonymous):

Do you mean: \[8\sqrt{63x^5} \ or\ 8(\sqrt{63x})^5\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which one do you want? The first or the second?

OpenStudy (anonymous):

the first

OpenStudy (anonymous):

\[8\sqrt{63x^5} = 8\sqrt{9*7x^5} = 8*3\sqrt{7x^5} = 24\sqrt{7x^5}\] \[=24\sqrt{7x*x^4} = ?\] What do you think the next step will be?

OpenStudy (anonymous):

making x^4 look like x*x*x*x

OpenStudy (anonymous):

Your goal is to get x^4 out of the square root? Can you do it as it is, or would it be easier to right x^4 as x^2*x^2?

OpenStudy (anonymous):

right x^4 as X^2*X^2

OpenStudy (anonymous):

So can you take something out of the square root now?

OpenStudy (anonymous):

we can take x^2 out

OpenStudy (anonymous):

Bravo! You can get it out as x^4 is x^2*x^2 = (x^2)^2!! So what would you have?

OpenStudy (anonymous):

24*2the sqrt of 9*7x

OpenStudy (anonymous):

\[24x^2\sqrt{7x}\] Do you want me to explain this?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

You can write \[x^2*x^2 \ as \ (x^2)^2\] Ok with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what would be the result of: \[\sqrt{(x^2)^2}\] ?

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

Let's try something easier. What is the result of: \[\sqrt{9}\]?

OpenStudy (anonymous):

The result will be a number, that when it's squared, it will produce the number 9. Which number is it?